OpenStudy (anonymous):

43 = 16 + 3x A. –19 B. 29 C. 9 D. 39

OpenStudy (nirmalnema):

43 = 16 + 3x 43-16=3x 27=3x x=27/3 x=9

OpenStudy (anonymous):

ok

OpenStudy (igreen):

@nirmalnema Please do not give answers, it's against OpenStudy's Code-of-Conduct. We are here to guide the asker to the answer, so we need his/her participation. http://openstudy.com/code-of-conduct

OpenStudy (nirmalnema):

lol ok..